The ages of Ashley and Mary are in the ratio 4:7. The sum of their ages is 22. Find the ages of Ashley and Mary.
Found 2 solutions by Maths68, lwsshak3:
Answer by Maths68(1474)

Let
Present age of Ashley =a
Present age of Mary =m
The ages of Ashley and Mary are in the ratio 4:7.
a:m=4:7
a/m=4/7
a=4m/7…………………….(1)
The sum of their ages is 22
a+m=22
a=22-m……………(2)
Put the value of a from (1) to (2)
a=22-m
4m/7=22-m
4m=7*(22-m)
4m=154-7m
4m+7m=154
11m=154
m=154/11
m=14
Put the value of m in (2)
a=22-m……………(2)
a=22-14
a=8
Present age of Ashley =a=8
Present age of Mary =m=14
Check
The ages of Ashley and Mary are in the ratio 4:7.
a:m=4:7
8:14=4:7
4:7=4:7
The sum of their ages is 22
a+m=22
8+14=22
22=22

Answer by lwsshak3(11628)

The ages of Ashley and Mary are in the ratio 4:7. The sum of their ages is 22. Find the ages of Ashley and Mary.
**
let x=Ashley's age
22-x=Mary's age
..
x/22-x=4/7
7x=88-4x
11x=88
x=8
ans:
Ashley's age=8
Mary's age=22-8=14